OpenStudy (anonymous):

What is the slope of the line 2x - 5y = 12?

OpenStudy (anonymous):

Is that a minecraft character?

OpenStudy (anonymous):

sorry off topic :P

OpenStudy (solomonzelman):

Put it into slope intercept form 2x - 5y = 12 - 5y = 12 - 2x y=5/2-12/5 Slope is 5/2

OpenStudy (solomonzelman):

I meant y=5/2x-12/5 Slope is 5/2

OpenStudy (anonymous):

Solve for y: -5y=-2x+12 Divide both sides by -5: y=(2/5)x + 12 Slope = 2/5

OpenStudy (ranga):

y = 2/5x - 12/5 slope = 2/5

OpenStudy (anonymous):

Sorry, I forgot to divide 12 by 5. Correct form would be: y = (2/5)x - 12/5

OpenStudy (anonymous):

Thanks all :D

OpenStudy (anonymous):

@victoria_lion Yeah, I have a youtube channel with 585 subscribers :P https://www.youtube.com/user/HelloWeAreGaming

OpenStudy (anonymous):

cool i'll subscribe